Hydrogeology

1. You are given the following data set from a falling head slug test. You are asked to determine the hydraulic conductivity of the aquifer using Hvorslevs method. Assume that the diameter of the well is 7.6 cm, the length of the screen is 98 cm and there is no gravel pack around the screen. Time (sec) head (m) h/ho 0.56 0.457 0.392 0.345 0.308 0.28 0.252 0.224 0.205 0.187 0.168 0.149 0.14 0.131 0.112 0.108 0.093 0.089 0.082 0.075 0.071 0.065 1.000 0.816 0.700 0.616 0.550 0.500 0.450 0.400 0.366 0.334 0.300 0.266 0.250 0.234 0.200 0.193 0.166 0.159 0.146 0.134 0.127 0.116 18 21 24 27 30 36 39 42 45 48 51 54 57 60 63
Answer #1

Solution:

In Horslev's Method:

hydraulic conductivity= k= \pi r2 * ln( h/h0) / F* TL

r= radius of screen

L=screen length

F=constant

This equation can be approximated by field data as

K = r2 *ln ( L/r) / 2*L* TL

TL = time lag value corresponding to the value h / h0 = 0.37

from the table h/h0 = 0.366 for t=24 = TL

ths, putting the values in the above equation:

K= (7.6/2)2 * ln( 98*2/ 7.6) / 2 * 98 * 24

=9.976 * 10-3 cm/s